# The art of modeling nuclear reactions with weakly bound nuclei: status and perspectives

### v1

@@ 9 Uncertainty evaluation
Given the amount of ingredients required for the computation of reaction observables (bound-state wavefuncions, optical potentials, spectroscopic factors,…) there has been a push for the quantification of the uncertainties in the calculated observables due to the ambiguities of these ingredients (the Igo ambiguity Igo58 of the optical potentials is a classical example of such ambiguities). While the classical response to this demand was the comparison of calculation using different ingredients (for example, optical potentials from different global parametrizations), this evaluation of the uncertainties is at most qualitative and highly dependent on the choice of inputs. Recently, some reaction theory groups, particularly the one at Michigan State University, have started to employ Bayesian methods to quantify the uncertainties that optical potentials suffer when they are obtained from the fitting of elastic scattering data (as is usual for optical potentials) and the propagation of these uncertainties to observables such as $(d,p)$ transfer angular or charge exchange $(p,n)$ differential cross sections Lov17b ; Lov18 ; Kin18 ; Kin19 ; Cat19 ; Cat20 ; Phi21 ; Whi22 or nucleon-knockout momentum distributions Heb23 . The determination of Bayesian posterior distributions and confidence intervals requires the evaluation of the reaction observables for a range of values for every considered parameter, which rapidly results in a humongous number of calculations, which are feasible in facilities with high computing power for light-computation models such as DWBA and ADWA, but become prohibitive for models which require heavier computations, such as CDCC or CRC. As such, these groups are exploring methods to accelerate these calculations using emulators Fur20 ; Dri21 to predict and interpolate the result of these heavier calculations for the required parameter values. Through these methods, uncertainty analyses have been performed using CDCC Sur22 , although admittedly only the uncertainties due to the valence-core interaction were studied, further proof of the computational magnitude of the task. In this regard, new global potential parametrizations with quantified uncertainties have recently been developed, with uncertainties originating either from the chiral interactions used to generate the potentials Whi21 or from fitting to extensive elastic scattering data Pru23 . The latter potential has been used to study the uncertainty in the reduction factors for transfer and knockout reactions Heb23b , which is still an open problem in Nuclear Physics Aum21 .
@@ 10 Inclusion of non-local potentials
It has long been known that nucleon-nucleus and nucleus-nucleus optical potentials are non-local Fes58a , both due to antisymmetrization and the effect of non-elastic channels (encoded in the second term of the Feshbach potential (3)). The use of non-local potentials transforms the Schrödinger equation into an integro-differential equation, which is far more cumbersome to solve, so local potentials have traditionally been favoured in reaction calculations. In fact, for a long time only a few non-local parametrizations for optical potentials have been published Per62 ; Gia76 ; Gia80 ; Tia15 and non-locality is not explicitly considered in the majority of potential parametrizations CH89 ; KD03 . In recent years there has been a development of dispersive optical potentials based on the nuclear self-energy, which is inherently non-local Cha07 ; Dic10 ; Mah16 ; Mor24 , thus reigniting the interest in fully non-local potentials.
For the study of elastic scattering, the neglect of non-locality is not problematic, as one can always find a local equivalent potential which reproduces the phase-shifts and observables of the non-local one Fie66 ; Fie67 . Analogously, reactions that are only sensitive to the asymptotics of the wave functions are usually well described by local potentials. However, non-locality reduces wavefunctions in the nuclear interior, in what is called the Perey effect Aus65 ; Fie66 , so non-locality should be taken into consideration in reactions that are sensitive to the nuclear interior. This has usually been approximated through the multiplication of wavefunctions obtained from a local potential by the so-called Perey factor Fie66 , which reduces the wavefunctions in the interior without altering their asymptotics. Although the Perey factor can be obtained for any non-locality shape, the prescriptions typically used assume a Gaussian dependence on non-locality Fra57 (usually referred to as Perey-Buck geometry). These approximations (local equivalent potential and Perey factor) have been compared to the computation of the wavefunction using the full non-local potential in Tit14 for $(d,p)$ transfer reactions, finding a difference in spectroscopic factors of $\sim 10\%$, although these results did not consider non-locality in the outgoing deuteron wavefunction.
@@ 11 Conclusions and perspectives
Some of the most promising advances are those related to few-body and ab-initio methods. The application of Faddeev method to nuclear reactions, while potentially providing an exact, rigorous solution of a given few-body Hamiltonian, has only been possible in recent years thanks to the increasing computational capabilities, the inclusion of complex optical potentials and the development of techniques to deal with the long-range Coulomb couplings. The method has been very useful for benchmarking some of the methods discussed in this review, such as the CDCC and ADWA methods Del07 ; Upa12 . It has also proven very useful for analyzing nucleon knockout reactions Cre11 ; Cre14 ; Cre16 .
@@
The expansion of the experimental studies, reaching heavier exotic nuclei, has also shown the need to extend and upgrade the existing reaction models to incorporate the new features found in the structure and reactions of the newly discovered nuclei. For example, it has become clear that a proper description of newly discovered halo nuclei in the islands of inversion requires the simultaneous inclusion of deformation, pairing and Pauli blocking. Reaction formalisms, such as DWBA or CDCC, must therefore be extended to accommodate such effects. Admittedly, the XCDCC method discussed in this review has used so far relatively simple particle-plus-rotor models, and hence it is mandatory to resort to more sophisticated models.
The challenge is even larger for the case of three-body systems which, in addition to the properties cited above, require specific formalisms to account for their three-body nature. For example, knockout experiments with 14Be have clear evidenced the need of including the excitations of the 12Be “core” Cor19 ; Cor23 . Even beyond, models including four- and five-body systems may be necessary, as the recent discovery of a tetraneutron correlated system in 8He attests Due22 .
All of these avenues for advancement show that, despite its numerous and significant successes, the study of nuclear reactions with weakly-bound nuclei remains a developing and active field, pushed by theoretical and technical advances and by new experimental measurements in ever more exotic nuclei. It is only fitting that an overview on such a field will be incomplete and become quickly outdated, but we hope we have managed to transmit to the reader at least a glimpse of the craft behind the study of such a vast and fascinating topic.
@@ Acknowledgements
The authors acknowledge financial support by MCIN/AEI/10.13039/501100011033 under I+D+i project No. PID2020-114687GB-I00. M.G.-R. acknowledges financial support by MCIN/AEI/10.13039/501100011033 under grant IJC2020-043878-I (also funded by “European Union NextGenerationEU/PRTR”).

### v2

@@ 10 Uncertainty evaluation
Given the amount of ingredients required for the computation of reaction observables (bound-state wavefuncions, optical potentials, spectroscopic factors,…) there has been a push for the quantification of the uncertainties in the calculated observables due to the ambiguities of these ingredients (the Igo ambiguity Igo58 of the optical potentials is a classical example of such ambiguities). While the classical response to this demand was the comparison of calculations using different ingredients (for example, optical potentials from different global parametrizations), this evaluation of the uncertainties is at most qualitative and highly dependent on the choice of inputs. Recently, Bayesian methods have been proposed to quantify the uncertainties that optical potentials suffer when they are obtained from the fitting of elastic scattering data (as is usual for optical potentials) and to study the propagation of these uncertainties to observables such as $(d,p)$ transfer angular or charge exchange $(p,n)$ differential cross sections Lov17b ; Lov18 ; Kin18 ; Kin19 ; Cat19 ; Cat20 ; Phi21 ; Whi22 or nucleon-knockout momentum distributions Heb23 . The determination of Bayesian posterior distributions and confidence intervals requires the evaluation of the reaction observables for a range of values of every considered parameter, which rapidly results in a humongous number of calculations, which are feasible in facilities with high computing power for light-computation models such as DWBA and ADWA, but become prohibitive for models which require heavier computations, such as CDCC or CRC. As such, methods to accelerate these calculations using emulators Fur20 ; Dri21 to predict and interpolate the result of these heavier calculations for the required parameter values are being explored. Through these methods, uncertainty analyses have been performed using CDCC Sur22 , although admittedly only the uncertainties due to the valence-core interaction were studied, further proof of the computational magnitude of the task. In this regard, new global potential parametrizations with quantified uncertainties have recently been developed, with uncertainties originating either from the chiral interactions used to generate the potentials Whi21 or from fitting to extensive elastic scattering data Pru23 , although it should be noted that in Pru23 the fit was not performed within a Bayesian framework. The latter potential has been used to study the uncertainty in the reduction factors for transfer and knockout reactions Heb23b , which is still an open problem Aum21 .
@@ 11 Inclusion of non-local potentials
It has long been known that nucleon-nucleus and nucleus-nucleus optical potentials are non-local Fes58a , both due to antisymmetrization and the effect of non-elastic channels (encoded in the second term of the Feshbach potential in Eq. (2)). The use of non-local potentials transforms the Schrödinger equation into an integro-differential equation, which is far more cumbersome to solve, so local potentials have traditionally been favoured in reaction calculations. In fact, for a long time only a few non-local parametrizations for optical potentials have been published Per62 ; Gia76 ; Gia80 ; Tia15 and non-locality is not explicitly considered in the majority of potential parametrizations CH89 ; KD03 . In recent years there has been a development of dispersive optical potentials based on the nuclear self-energy, which is inherently non-local Cha07 ; Dic10 ; Mah16 ; Mor24 , thus reigniting the interest in fully non-local potentials.
For the study of elastic scattering, the neglect of non-locality is not problematic, as one can always find a local equivalent potential which reproduces the phase-shifts and observables of the non-local one Fie66 ; Fie67 . Analogously, reactions that are only sensitive to the asymptotics of the wave functions are usually well described by local potentials. However, non-locality reduces wavefunctions in the nuclear interior, in what is called the Perey effect Aus65 ; Fie66 , so non-locality should be taken into consideration in reactions that are sensitive to the nuclear interior, such as transfer reactions. This has usually been approximated through the multiplication of wavefunctions obtained from a local potential by the so-called Perey factor Fie66 , which reduces the wavefunctions in the interior without altering their asymptotics. Although the Perey factor can be obtained for any non-locality shape, the prescriptions typically used assume a Gaussian dependence on non-locality Fra57 (usually referred to as Perey-Buck geometry). These approximations (local equivalent potential and Perey factor) have been compared to the computation of the wavefunction using the full non-local potential in Tit14 for $(d,p)$ transfer reactions, finding a difference in spectroscopic factors of $\sim 10\%$, although these results did not consider non-locality in the outgoing deuteron wavefunction.
@@ 12 Conclusions and perspectives
  • Some of the most promising advances are those related to few-body and ab-initio methods. The application of the Faddeev method to nuclear reactions, while potentially providing an exact, rigorous solution of a given few-body Hamiltonian, has only been possible in recent years thanks to the increasing computational capabilities, the inclusion of complex optical potentials and the development of techniques to deal with the long-range Coulomb couplings. The method has been very useful for benchmarking some of the methods discussed in this review, such as the CDCC and ADWA methods Del07 ; Upa12 . It has also proven very useful for analyzing nucleon knockout reactions Cre11 ; Cre14 ; Cre16 .
@@
  • The expansion of experimental studies, reaching heavier exotic nuclei, has also shown the need to extend and upgrade the existing reaction models to incorporate the new features found in the structure and reactions of the newly discovered nuclei. For example, it has become clear that a proper description of newly discovered halo nuclei in the islands of inversion requires the simultaneous inclusion of deformation, pairing and Pauli blocking. Reaction formalisms, such as DWBA or CDCC, must therefore be extended to accommodate such effects. Admittedly, the XCDCC method discussed in this review has used so far relatively simple particle-plus-rotor models, and hence it is mandatory to resort to more sophisticated models.
  • The challenge is even larger for the case of three-body systems which, in addition to the properties cited above, require specific formalisms to account for their three-body nature. For example, knockout experiments with 14Be have clear evidenced the need of including the excitations of the 12Be “core” Cor19 ; Cor23 . Even beyond, models including four- and five-body systems may be necessary, as the recent discovery of a four-neutron correlated system in 8He attests Due22 .
All of these avenues for advancement show that, despite its numerous and significant successes, the study of nuclear reactions with weakly-bound nuclei remains a developing and active field, pushed by theoretical and technical advances and by new experimental measurements in ever more exotic nuclei. While we hope that this work has transmitted to the reader at least a glimpse of the current methods in this fascinating field, we also hold the paradoxical hope that new developments promptly render this work somewhat obsolete thanks to new and perhaps surprising advancements.
@@ Acknowledgements
We are grateful to Angela Bonaccorso and Jin Lei for their feedback on the semiclassical transfer to the continuum calculations. This work has been partially funded by the Ministerio de Ciencia e Innovación, MCIN/AEI/10.13039/501100011033 under I+D+i project No. PID2020-114687GB-I00. M.G.-R. acknowledges financial support by MCIN/AEI/10.13039/501100011033 under grant IJC2020-043878-I (also funded by “European Union NextGenerationEU/PRTR”).